# Reasoning-Enhanced Domain-Adaptive Pretraining of Multimodal Large Language Models for Short Video Content Moderation

### v1

@@ Abstract
Reasoning-Enhanced Domain-Adaptive Pretraining of Multimodal Large Language Models for Short Video Content Moderation
Zixuan Wangthanks: Equal contribution., Yu Sun11footnotemark: 1 Hongwei Wang, Baoyu Jing, Xiang Shen, Xin Dong, Zhuolin Hao Hongyu Xiong, Yang Song TikTok
@@ 1 Introduction
Figure 1: Illustration of our domain-adaptive pretraining approach for short video content moderation. For each issue type, we first decompose the annotation guidelines into a set of sub-questions to assist pretraining data generation. An annotator MLLM is deployed to produce three types of pretraining data: Caption, VQA, and CoT, enabling the model to mimic human-like reasoning. The model can be pretrained using three different strategies, and the pretrained model is finally evaluated in both zero-shot and SFT settings.
@@
Short video platforms such as Reels and YouTube Shorts have experienced explosive growth. While these platforms offer unprecedented opportunities for communication and information dissemination, they have simultaneously facilitated the distribution of a wide range of inappropriate content. Videos featuring borderline sexual content, plagiarism, and fake engagement, not only affect user experience but also deteriorates content ecosystem in the long run. As a result, the regulation and moderation of short video content have become increasingly critical to ensure a healthy digital ecosystem.
@@
Current short video platforms typically implement a standardized content moderation process, which consists of three sequential stages: The platform initially develops comprehensive guidelines for identifying specific issues (such as sexually suggestive content); Subsequently, human annotators are instructed to label video samples according to these guidelines; Finally, this labeled dataset serves as labeled data to fine-tune visual models (e.g., VLMo Bao et al. (2022b), BEiT Bao et al. (2022a), and X-VLM Wang et al. (2022)) for video classification, which are then deployed for real-time content monitoring and filtering. However, this process comes with significant limitations: (1) High manual labeling cost. These visual models are typically small in size (usually only a few hundred megabytes) and lack world knowledge, so they require large volumes of manually labeled data, making the process time-consuming and expensive. (2) No cross-issue generalization. Each of these small models is designed to handle only a fixed issue and lacks the ability to generalize across different issues. Therefore, when the definition of an issue changes or a new issue emerges, the model must be retrained from scratch following the same labor-intensive process, leading to long development cycles.
@@
To address the limitations of small specialized models, we explore the use of multi-modal large language models (MLLMs). MLLMs possess extensive world knowledge, which theoretically reduces the need for large-scale annotated data and offers better cross-issue generalization. However, applying MLLMs in this context presents two key challenges: (1) Domain-specific data distribution. The data on short video platforms exhibits unique characteristics, both in format (e.g., frame layouts, visual composition, and hashtag usage patterns) and in semantics (e.g., narrative styles, editing conventions, and internet slang). These characteristics often diverge significantly from the distribution of data MLLMs were originally pretrained on, potentially impacting performance. (2) Complex and evolving definition of issues. Each moderation issue comes with its own well-defined criteria and decision rules, which are usually complicated and may evolve. For an MLLM to make accurate judgments, it must thoroughly understand and follow these guidelines to perform reliable inference.
@@
To tackle the above challenges, we propose a novel domain-adaptive pretraining approach for short video moderation. As illustrated in Figure 1, the pretraining approach includes three key components: (1) Enhanced video detail perception. To help MLLMs better adapt to the unique data patterns of short videos, we introduce a descriptive video captioning (Caption) task into the pretraining stage. These captions guide the model to attend to fine-grained video detail, especially those relevant to specific issues. (2) Deep understanding of guidelines. To enable MLLMs to thoroughly understand the nuances of issue annotation guidelines, we decompose each guideline into a set of sub-questions and construct a visual question answering (VQA) task. For example, in the case of borderline sexual content, the guideline is broken down into questions such as “Are private body parts exposed?”, with annotated answers. This enables MLLMs to learn the specific judgment criteria for each issue. (3) Structured reasoning. Even after perceiving video details and understanding the guidelines, the model must follow a systematic reasoning process. To support this, we curate a chain-of-thought (CoT) task that teaches MLLMs to reason step by step following predefined logic flows. Together, the Caption, VQA, and CoT datasets provide end-to-end domain adaptation for MLLMs, which significantly improves the performance.
@@
We conducted the pretraining on several open-source MLLMs, including LLaVA-OV Li et al. (2024) and Qwen2.5-VL Bai et al. (2024), using data from three content moderation tasks: sexually suggestive content, unoriginal content, and fake engagement. Experimental results show that our domain-adaptive pretraining significantly improves model performance compared with native models, both in zero-shot and supervised fine-tuning (SFT) settings. Moreover, we demonstrate that our approach enables strong generalization to unseen issues with minimal supervision, highlighting its robustness and adaptivity. Our ablation studies further validate the effectiveness of our method across different model scales.
@@
  • We propose an MLLM pretraining framework for short video moderation, demonstrating strong generalization to out-of-distribution issues not seen during pretraining.
@@ 2 Related Work
Video Content Moderation. Recent advances in deep neural networks have significantly enhanced content moderation capabilities in detecting and classifying inappropriate video content on short video platforms Yousaf and Nawaz (2022); Tanantong and Yongwattana (2023). For example, multimodal features have been applied to identify deceptive clickbait videos Rahman et al. (2023); Sun et al. (2025); Liu et al. (2024a); Liang et al. (2025); age-adaptive learning techniques have been developed to detect and categorize video content inappropriate for different viewer demographics, demonstrating sensitivity to age-related moderation needs Alam et al. (2024). These approaches underscore the potential of MLLMs in addressing the challenges of short video content moderation.
Domain-Adaptive Pretraining. Recent research demonstrates the effectiveness of domain-adaptive pretraining in enhancing model performance Song et al. (2025). Specifically, pretraining strategies have been used in model design, task formulation, and application scenarios Guo et al. (2024). For example, in multimodal learning, efficient domain-specific pretraining has been applied to human activity recognition Bulat et al. (2024) and short video understanding Lu et al. (2025); SimRAG Hong et al. (2024); Xu et al. (2024) and domain-specific instruction tuning Xie et al. (2024); Ke et al. (2023); Roth et al. (2024); Cheng et al. (2024); Wang et al. (2025); Dong et al. (2025) effectively align models with VQA tasks Ging et al. (2024); Khullar et al. (2024); adaptive techniques such as velocity-based domain reweighting Luo et al. (2024) and structure-aware knowledge injection Liu et al. (2024b) further improve knowledge retention and transfer.
@@ 3 The Proposed Pretraining Approach
3.1 A Unified Model for All Moderation Issues
Short video platforms often face a variety of content moderation issues, such as non-original content. We denote the set of all issues of interest by $\mathcal{I}$. For each issue $i\in\mathcal{I}$, the moderation team defines detailed descriptions and corresponding annotation guidelines, denoted by $G_{i}$. Human annotators are then trained using these guidelines to label a collection of videos, resulting in a dataset $\mathcal{D}_{i}$. This dataset is then used to fine-tune a visual classification model $\mathcal{M}_{i}$:
@@ 3.3 Pretraining Tasks
There are two main challenges in using MLLMs for short video content moderation: (1) The unique characteristics of short videos limit the capabilities of native MLLMs to fully understand their content; (2) Issue annotation guidelines are often complex, making it difficult for native MLLMs to effectively follow and reason based on them. To address these challenges, we design three pretraining tasks:
@@ 5.4 Case Study
We conducted a case study to intuitively demonstrate the effect of pretraining. As shown in Figure 4 in the Appendix, the Spanish text in the video frame contains an obvious sexually suggestive message. However, the model without pretraining fails to recognize the meaning of the text and subsequently produces incorrect reasoning. In contrast, the pretrained model provides both an accurate answer and a correct reasoning process.
@@ 6 Conclusion
In this work, we propose a domain-adaptive pretraining approach for short video content moderation. We design three novel pretraining tasks: Caption, VQA, and CoT, which effectively enhance MLLMs’ ability to perceive video details, understand annotation guidelines, and perform reasoning. Experimental results show that our pretraining method significantly improves the model’s performance on both zero-shot and SFT evaluations. Furthermore, the pretrained models exhibit strong generalization to new issues, which can substantially reduce manual annotation costs and shorten development cycles in real-world deployment scenarios.
@@ 7 Limitation
In this work, the largest model we used is 7B due to limitations in resources. Although the pretrained 7B models already perform well for short video content moderation, we are still interested in exploring the impact of pretraining on even larger models. Regarding data types used for pretraining, in addition to MLLM-annotated data, we also plan to incorporate a large amount of unlabeled data to further enhance performance. Moreover, when decomposing annotation guidelines into sub-questions, we currently rely on manual design, which may not be optimal. In the future, we plan to explore reinforcement learning methods to automatically decompose annotation guidelines in a more efficient and adaptive manner.
@@ Appendix A Details of Prompt Design
Appendix B Details of Case Study
In Figure 4, we present a detailed example [^1] This example is for illustrative purposes only and does not depict a real post. that violates the SSC policy, along with the MLLM’s responses before and after pretraining.
Appendix C Analysis on Qwen’s Inconsistency Issue

### v2

# Reasoning-Enhanced Domain-Adaptive Pretraining of Multimodal Large Language Models for Short Video Content Governance

@@ Abstract
Reasoning-Enhanced Domain-Adaptive Pretraining of Multimodal Large Language Models for Short Video Content Governance
Zixuan Wangthanks: Equal contribution., Yu Sun11footnotemark: 1, Hongwei Wang, Baoyu Jing, Xiang Shen, Xin Dong, Zhuolin Hao, Hongyu Xiong, Yang Song TikTok Inc. {zixuan.wang1, yu.sun, hongwei.w, baoyu.jing, xindong, haozhuolin, hongyu.xiong}@tiktok.com
@@ 1 Introduction
Figure 1: Illustration of our domain-adaptive pretraining approach for short video content governance. For each issue type, we first decompose the annotation guidelines into a set of sub-questions to assist pretraining data generation. An annotator MLLM is deployed to produce three types of pretraining data: Caption, VQA, and CoT, enabling the model to mimic human-like reasoning. The model can be pretrained using three different strategies, and the pretrained model is finally evaluated in both zero-shot and SFT settings.
@@
Short video platforms such as Reels and YouTube Shorts have experienced explosive growth. While these platforms offer unprecedented opportunities for communication and information dissemination, they have simultaneously facilitated the distribution of a wide range of inappropriate content. Videos featuring borderline sexual content, plagiarism, and fake engagement, not only affect user experience but also deteriorates content ecosystem in the long run. As a result, the regulation and governance of short video content have become increasingly critical to ensure a healthy digital ecosystem.
@@
Current short video platforms typically implement a standardized content governance process, which consists of three sequential stages: The platform initially develops comprehensive guidelines for identifying specific issues (such as sexually suggestive content); Subsequently, human annotators are instructed to label video samples according to these guidelines; Finally, this labeled dataset serves as labeled data to fine-tune visual models (e.g., VLMo Bao et al. (2022b), BEiT Bao et al. (2022a), and X-VLM Wang et al. (2022)) for video classification, which are then deployed for real-time content monitoring and filtering. However, this process comes with significant limitations: (1) High manual labeling cost. These visual models are typically small in size (usually only a few hundred megabytes) and lack world knowledge, so they require large volumes of manually labeled data, making the process time-consuming and expensive. (2) No cross-issue generalization. Each of these small models is designed to handle only a fixed issue and lacks the ability to generalize across different issues. Therefore, when the definition of an issue changes or a new issue emerges, the model must be retrained from scratch following the same labor-intensive process, leading to long development cycles.
@@
To address the limitations of small specialized models, we explore the use of multi-modal large language models (MLLMs). MLLMs possess extensive world knowledge, which theoretically reduces the need for large-scale annotated data and offers better cross-issue generalization. However, applying MLLMs in this context presents two key challenges: (1) Domain-specific data distribution. The data on short video platforms exhibits unique characteristics, both in format (e.g., frame layouts, visual composition, and hashtag usage patterns) and in semantics (e.g., narrative styles, editing conventions, and internet slang). These characteristics often diverge significantly from the distribution of data MLLMs were originally pretrained on, potentially impacting performance. (2) Complex and evolving definition of issues. Each content issue comes with its own well-defined criteria and decision rules, which are usually complicated and may evolve. For an MLLM to make accurate judgments, it must thoroughly understand and follow these guidelines to perform reliable inference.
@@
To tackle the above challenges, we propose a novel domain-adaptive pretraining approach for short video governance. As illustrated in Figure 1, the pretraining approach includes three key components: (1) Enhanced video detail perception. To help MLLMs better adapt to the unique data patterns of short videos, we introduce a descriptive video captioning (Caption) task into the pretraining stage. These captions guide the model to attend to fine-grained video detail, especially those relevant to specific issues. (2) Deep understanding of guidelines. To enable MLLMs to thoroughly understand the nuances of issue annotation guidelines, we decompose each guideline into a set of sub-questions and construct a visual question answering (VQA) task. For example, in the case of borderline sexual content, the guideline is broken down into questions such as “Are private body parts exposed?”, with annotated answers. This enables MLLMs to learn the specific judgment criteria for each issue. (3) Structured reasoning. Even after perceiving video details and understanding the guidelines, the model must follow a systematic reasoning process. To support this, we curate a chain-of-thought (CoT) task that teaches MLLMs to reason step by step following predefined logic flows. Together, the Caption, VQA, and CoT datasets provide end-to-end domain adaptation for MLLMs, which significantly improves the performance.
@@
We conducted the pretraining on several open-source MLLMs, including LLaVA-OV Li et al. (2024) and Qwen2.5-VL Bai et al. (2024), using data from three content governance tasks: sexually suggestive content, unoriginal content, and fake engagement. Experimental results show that our domain-adaptive pretraining significantly improves model performance compared with native models, both in zero-shot and supervised fine-tuning (SFT) settings. Moreover, we demonstrate that our approach enables strong generalization to unseen issues with minimal supervision, highlighting its robustness and adaptivity. Our ablation studies further validate the effectiveness of our method across different model scales.
@@
  • We propose an MLLM pretraining framework for short video governance, demonstrating strong generalization to out-of-distribution issues not seen during pretraining.
@@ 2 Related Work
2.1 Video Content Governance
Recent advances in deep neural networks have significantly enhanced content governance capabilities in detecting and classifying inappropriate video content on short video platforms Yousaf and Nawaz (2022); Li et al. (2025b). For example, multimodal features have been applied to identify misleading clickbait videos Rahman et al. (2023); Sun et al. (2025); Liang et al. (2025); Age-adaptive learning techniques have been developed to detect and categorize inappropriate video content for different demographics of viewers, demonstrating sensitivity to visual needs Alam et al. (2024); Zhang et al. (2024); Li et al. (2025a). These approaches underscore the potential of MLLMs in addressing the challenges of short video content governance.
2.2 Domain-Adaptive Pretraining
Recent research demonstrates the effectiveness of domain-adaptive pretraining in enhancing model performance Song et al. (2025). Specifically, pretraining strategies have been used in model design, task formulation, and application scenarios Guo et al. (2024). For example, in multimodal learning, efficient domain-specific pretraining has been applied to human activity recognition Bulat et al. (2024) and short video understanding Lu et al. (2025); Li et al. (2023); SimRAG Hong et al. (2024) and domain-specific instruction tuning Xie et al. (2024); Liu et al. (2024a); Wang et al. (2025); Dong et al. (2025) effectively align models with VQA tasks Ging et al. (2024); Khullar et al. (2024); Adaptive techniques such as velocity-based domain reweighting Luo et al. (2024); Leong et al. (2025), knowledge distillation Shi et al. (2024) and structure-aware knowledge injection Liu et al. (2024b) further improve knowledge retention and transfer.
@@ 3 The Proposed Pretraining Approach
3.1 A Unified Model for All Content Issues
Short video platforms often face a variety of content issues, such as non-original content. We denote the set of all issues of interest by $\mathcal{I}$. For each issue $i\in\mathcal{I}$, the policy team defines detailed descriptions and corresponding annotation guidelines, denoted by $G_{i}$. Human annotators are then trained using these guidelines to label a collection of videos, resulting in a dataset $\mathcal{D}_{i}$. This dataset is then used to fine-tune a visual classification model $\mathcal{M}_{i}$:
@@ 3.3 Pretraining Tasks
There are two main challenges in using MLLMs for short video content governance: (1) The unique characteristics of short videos limit the capabilities of native MLLMs to fully understand their content; (2) Issue annotation guidelines are often complex, making it difficult for native MLLMs to effectively follow and reason based on them. To address these challenges, we design three pretraining tasks:
@@ 5.4 Case Study
We conducted a case study to intuitively demonstrate the effect of pretraining. As shown in Figure 3, the Spanish text in the video frame contains an obvious sexually suggestive message. However, the model without pretraining fails to recognize the meaning of the text and subsequently produces incorrect reasoning. In contrast, the pretrained model provides both an accurate answer and a correct reasoning process.
@@ 6 Conclusion
In this work, we propose a domain-adaptive pretraining approach for short video content governance. We design three novel pretraining tasks: Caption, VQA, and CoT, which effectively enhance MLLMs’ ability to perceive video details, understand annotation guidelines, and perform reasoning. Experimental results show that our pretraining method significantly improves the model’s performance on both zero-shot and SFT evaluations. Furthermore, the pretrained models exhibit strong generalization to new issues, which can substantially reduce manual annotation costs and shorten development cycles in real-world deployment scenarios.
@@ 7 Limitation
In this work, the largest model we used is 7B due to limitations in resources. Although the pretrained 7B models already perform well for short video content governance, we are still interested in exploring the impact of pretraining on even larger models. Regarding data types used for pretraining, in addition to MLLM-annotated data, we also plan to incorporate a large amount of unlabeled data to further enhance performance. Moreover, when decomposing annotation guidelines into sub-questions, we currently rely on manual design, which may not be optimal. In the future, we plan to explore reinforcement learning methods to automatically decompose annotation guidelines in a more efficient and adaptive manner.
@@ Appendix A Details of Prompt Design
Appendix B Analysis on Qwen’s Inconsistency Issue